The gravitational force exerted on a solid object is 113 N. when the object is suspended from a spring scale and submerged in water (Density of water=1000 kg/m3), the scale reads 50 N. The density of this object (in kg/m3) is:


1
Expert's answer
2021-11-25T10:14:24-0500

"\frac{\rho-1000}{\rho}=\frac{50}{113}\\\rho=1794\frac{kg}{m^3}"